A genetic test is use to determine if people have a disposition for thrombosis, which is the formation of a blood clot inside a blood vessel that obstructs the flow of blood though the circulatory system. It is believed that 3% of people actually have this predisposition. The genetic test is 99% accurate if a person has the predisposition, meaning that the probability of a positive test result when a person actually has the predisposition is 0.99. The test is 98% accurate if a person does not have the predisposition. What is the probability that a randomly selected person who test positive for the predisposition by the test actually has the predisposition?
Answer #1
P(tested positive)=P(has predisposition)*P(tested positive|has predisposition)+P(not has predisposition)*P(tested positive|not has predisposition)
=0.03*0.99+0.97*(1-0.98)=0.0491
therefore P(has predisposition|tested positive)=P(has predisposition)*P(tested positive|has predisposition)/P(tested positive)=0.03*0.99/0.0491 =0.6049